Bicycling World, a magazine devoted to cycling, reviews hundreds of bicycles throughout the year. Its "Road-Race" category contains reviews of bicycles used by riders primarily interested in racing. One of the most important factors in selecting a bicycle for racing is the weight of the bicycle. The following data show the weight (pounds) and price ($) for ten racing bicycles reviewed by the magazine

Model Weight (lb) Price ($)
  Fierro 7B 17.9 2,200
  HX 5000 16.2 6,350
  Durbin Ultralight 15.0 8,470
  Schmidt 16.0 6,300
  WSilton Advanced 17.3 4,100
  Bicyclette vélo 13.2 8,700
  Supremo Team 16.3 6,100
  XTC Racer 17.2 2,680
  D'Onofrio Pro 17.7 3,500
  Americana #6 14.2 8,100


Regression Model
Y = 28818.00 + -1439.01

The manufacturers of the D'Onofrio Pro plan to introduce the 13.6 lb D'Onofrio Elite bicycle later this year. Use the regression model you estimated in part (b) to predict the price of the D'Ononfrio Elite.
Answer #1

Answer

Given regression equation is

y= 28818 - 1439.01(weight)

setting weight = 13.6

we will get

y= 28818 -1439.01(13.6)

= 28818 - 19570.54

= 9247.46